# Optimal Computation in Anonymous Dynamic Networks

### v7

@@ 2.1 Model of Computation
Processes and networks. A dynamic network on a finite non-empty set $V=\{p_{1},p_{2},\dots,p_{n}\}$ is an infinite sequence $\mathcal{G}=(G_{t})_{t\geq 1}$, where $G_{t}=(V,E_{t})$ is an undirected multigraph, i.e., $E_{t}$ is a multiset of unordered pairs of elements of $V$. In this context, the set $V$ is called system, and its $n\geq 1$ elements are the processes. The elements of the multiset $E_{t}$ are called links; note that we allow any (finite) number of “parallel links” between two processes, as well as “self-loops”. [^11] In the dynamic networks literature, $G_{t}$ is typically assumed to be a simple graph, for at most one link between the same two processes is allowed. However, our results hold more generally for multigraphs.
@@
Input and internal states. Each process $p_{i}$ starts with an input $\lambda(p_{i})$, which is assigned to it at round 00. It also has an internal state, which is initially determined by $\lambda(p_{i})$. At each round $t\geq 1$, every process composes a message (depending on its internal state) and broadcasts it to its neighbors in $G_{t}$ through all its incident links. [^12] In order to model wireless radio communication, it is natural to assume that each process in a dynamic network broadcasts its messages to all its neighbors (a message is received by anyone within communication range). The network’s anonymity prevents processes from specifying single destinations. By the end of round $t$, each process reads all messages coming from its neighbors and updates its internal state according to a local algorithm $\mathcal{A}$. That is, $\mathcal{A}$ takes as input the current internal state of the process and the multiset of incoming messages, and returns the new internal state of the process. Note that $\mathcal{A}$ is deterministic and is the same for all processes. [^13] This network model bears some similarities with Population Protocols, although there are radical differences in the way symmetry is handled. The point-to-point communication model of Population Protocols automatically breaks the symmetry between communicating agents, greatly simplifying problems such as Leader Election, Average Consensus, etc. The input of each process also includes a leader flag. The processes whose leader flag is set are called leaders (or supervisors). We will denote the number of leaders by $\ell$.
@@
Stabilization and termination. Each process returns an output at the end of each round, which is determined by its current internal state. A system is said to stabilize if the outputs of all its processes remain constant from a certain round onward; note that a process’ internal state may still change even when its output is constant. A process may also decide to explicitly terminate and no longer update its internal state. When all processes have terminated, the system is said to terminate, as well.
@@
Computation. We say that an algorithm $\mathcal{A}$ computes a function $F$ if, whenever the processes in the system are assigned inputs $\lambda(p_{1})$, $\lambda(p_{2})$, …, $\lambda(p_{n})$ and all processes execute the local algorithm $\mathcal{A}$ at every round, the system eventually stabilizes with each process $p_{i}$ giving the desired output $F(p_{i},\lambda)$. [^14] Formally, a function computed by a system of $n$ processes maps $n$-tuples of input values to $n$-tuples of output values. Writing such a function as $F(p_{i},\lambda)$ emphasizes that the output of a process may depend on all processes’ inputs, as well as on the process itself. That is, different processes may give different outputs. A stronger notion of computation requires the system to not only stabilize but also to explicitly terminate with the correct output. The (worst-case) running time of $\mathcal{A}$, as a function of $n$, is the maximum number of rounds it takes for the system to stabilize (and optionally terminate), taken across all possible dynamic networks of $n$ processes and all possible input assignments.
@@ 2.2 Classes of Functions
Multiset-based functions. Let $\mu_{\lambda}=\{(z_{1},m_{1}),(z_{2},m_{2}),\dots,(z_{k},m_{k})\}$ be the multiset of all processes’ inputs. That is, for all $1\leq i\leq k$, there are exactly $m_{i}$ processes $p_{j_{1}},p_{j_{2}},\dots,p_{j_{m_{i}}}$whose input is $z_{i}=\lambda(p_{j_{1}})=\lambda(p_{j_{2}})=\dots=\lambda(p_{j_{m_{i}}})$; note that $n=\sum_{i=1}^{k}m_{i}$. A function $F$ is multiset-based if it has the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}),\mu_{\lambda})$. That is, the output of each process depends only on its own input and the multiset of all processes’ inputs. [^15] Multiset-based functions are also know as multi-aggregation or multi-aggregate functions [23, 25].
@@
Once a process $p_{i}$ with input $\lambda(p_{i})$ has determined the multiset $\mu_{\lambda}$ of all processes’ inputs, it can immediately compute any desired function $\psi(\lambda(p_{i}),\mu_{\lambda})$ within the same local computation phase. ∎
@@
The problem of computing the mean of all input values is called Average Consensus  [7, 16, 26, 34, 44, 45, 48, 49, 50, 55, 58]. The frequency-based function $F_{IF}(p_{i},\lambda)=\frac{1}{n}\cdot\mu_{\lambda}$ is called Input Frequency function, and is complete for the class of frequency-based functions.
@@
Suppose that a process $p_{i}$ has determined $\frac{1}{n}\cdot\mu_{\lambda}=\{(z_{1},m_{1}/n),(z_{2},m_{2}/n),\dots,(z_{k},m_{k}/n)\}$. Then it can immediately find the smallest integer $d>0$ such that $d\cdot(m_{i}/n)$ is an integer for all $1\leq i\leq k$. Note that $\frac{d}{n}\cdot\mu_{\lambda}$ is a multiset. Hence, in the same round, $p_{i}$ can compute any desired function $\psi(\lambda(p_{i}),\frac{d}{n}\cdot\mu_{\lambda})$, and thus any frequency-based function, by definition. ∎
@@ 2.3 Disconnected Networks
Although the network $G_{t}$ at each individual round may be disconnected, in this paper we assume dynamic networks to be $\tau$-union-connected. That is, there is a (smallest) parameter $\tau\geq 1$, called dynamic disconnectivity, such that the sum of any $\tau$ consecutive $G_{t}$’s is a connected multigraph. [^16] By definition, the sum of (multi)graphs is obtained by adding together their adjacency matrices. Thus, for all $i\geq 1$, the multigraph $\left(V,\bigcup_{t=i}^{i+\tau-1}E_{t}\right)$ is connected (we remark that a union of multisets adds together the multiplicities of equal elements). [^17] Our $\tau$-union-connected networks should not be confused with the $\tau$-interval-connected networks from [38]. In those networks, the intersection (as opposed to the union) of any $\tau$ consecutive $E_{t}$’s induces a connected (multi)graph. In particular, a $\tau$-interval-connected network is connected at every round, while a $\tau$-union-connected network may not be, unless $\tau=1$. Incidentally, a network is 1-interval-connected if and only if it is 1-union-connected.
We will prove that terminating computations are impossible without some a-priori knowledge about $\tau$. We say that a function $F(p_{i},\lambda)$ is trivial if and only if it is of the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}))$. That is, the output of any process $p_{i}$ depends only on its own input $\lambda(p_{i})$ and not on the inputs of other processes; therefore, $F$ can be computed “locally” and does not require communications between processes.
@@
Any non-trivial function is impossible to compute with termination unless the processes have some knowledge about $\tau$.
@@
Assume for a contradiction that a non-trivial function $F(p_{i},\lambda)$ is computed with termination by an algorithm $\mathcal{A}$ with no knowledge of $\tau$. Since $F$ is non-trivial, there is an input value $z$ and two distinct output values $y$ and $y^{\prime}$ with the following properties. (i) If $p_{1}$ is the only process in a network (i.e., $n=1$), and $p_{1}$ is assigned input $z$ and executes $\mathcal{A}$, it terminates in $t$ rounds with output $y$. (ii) There exists a network size $\tilde{n}>1$ and an input assignment $\lambda$ with $\lambda(p_{1})=z$ such that, whenever the processes in a network of size $\tilde{n}$ are assigned input $\lambda$ and execute $\mathcal{A}$, the process $p_{1}$ eventually terminates with output $y^{\prime}\neq y$.
Let us now consider a dynamic network of $\tilde{n}$ processes, where $p_{1}$ is kept disconnected from all other processes for the first $t$ rounds (hence $\tau>t$). Assign input $\lambda$ to the processes, and let them execute algorithm $\mathcal{A}$. Due to property (i), since $p_{1}$ is isolated for $t$ rounds and has no knowledge of $\tau$, it terminates in $t$ rounds with output $y$. This contradicts property (ii), which states that $p_{1}$ should terminate with output $y^{\prime}\neq y$. ∎
The following result will be used repeatedly to reduce computations on $\tau$-union connected networks to computations on 1-union-connected networks.
@@
A function $F$ can be computed (with termination) within $f(n)$ rounds in any dynamic network with $\tau=1$ if and only if $F$ can be computed (with termination) within $\tau\cdot f(n)$ rounds in any dynamic network with $\tau\geq 1$, assuming that $\tau$ is known to all processes.
@@
Subdivide time into blocks of $\tau$ consecutive rounds, and consider the following algorithm. Each process collects and stores all messages it receives within a same block, and updates its state all at once at the end of the block. This reduces any $\tau$-union-connected network $\mathcal{G}=((V,E_{t}))_{t\geq 1}$ to a 1-union-connected network $\mathcal{G}^{\prime}=((V,E^{\prime}_{t}))_{t\geq 1}$, where $E^{\prime}_{t}=\bigcup_{i=(t-1)\tau+1}^{t\tau}E_{i}$. Thus, if $F$ can be computed within $f(n)$ rounds in all 1-union-connected networks (which include $\mathcal{G}^{\prime}$), then $F$ can be computed within $\tau f(n)$ rounds in the original network $\mathcal{G}$ [^18] Note that this argument is correct because algorithms are required to work for all multigraphs, as opposed to simple graphs only. Indeed, since a process $p_{i}$ may receive multiple messages from the same process $p_{j}$ within a same block, the resulting network $\mathcal{G}^{\prime}$ may have multiple links between $p_{i}$ and $p_{j}$ in a same round, even if $\mathcal{G}$ does not.
@@
Conversely, consider a 1-union-connected network $\mathcal{G}$, and construct a $\tau$-union-connected network $\mathcal{G}^{\prime}$ by inserting $\tau-1$ empty rounds (i.e., rounds with no links at all) between every two consecutive rounds of $\mathcal{G}$. Since no information circulates during the empty rounds, if $F$ cannot be computed within $f(n)$ rounds in $\mathcal{G}$, then $F$ cannot be computed within $\tau f(n)$ rounds in $\mathcal{G}^{\prime}$ (recall that running times are measured in the worst case across all possible networks). ∎
@@
Relationship between dynamic disconnectivity and the dynamic diameter. A concept closely related to the dynamic disconnectivity $\tau$ of a network is its dynamic diameter (or temporal diameter) $d$, which is defined as the maximum number of rounds it may take for information to travel from any process to any other process at any point in time [11, 40]. It is a simple observation that $\tau\leq d\leq\tau(n-1)$.

### v8

@@ 2.1 Model of Computation
Agents and networks. A dynamic network on a finite non-empty set $V=\{p_{1},p_{2},\dots,p_{n}\}$ is an infinite sequence $\mathcal{G}=(G_{t})_{t\geq 1}$, where $G_{t}=(V,E_{t})$ is an undirected multigraph, i.e., $E_{t}$ is a multiset of unordered pairs of elements of $V$. In this context, the set $V$ is called system, and its $n\geq 1$ elements are the agents. The elements of the multiset $E_{t}$ are called links; note that we allow any (finite) number of “parallel links” between two agents, as well as “self-loops”. Note that, in the dynamic networks literature, $G_{t}$ is typically assumed to be a simple graph, for at most one link between the same two agents is allowed. However, our results hold more generally for multigraphs.
@@
Input and internal states. Each agent $p_{i}$ starts with an input $\lambda(p_{i})$, which is assigned to it at round 00. It also has an internal state, which is initially determined by $\lambda(p_{i})$. At each round $t\geq 1$, every agent composes a message (depending on its internal state) and broadcasts it to its neighbors in $G_{t}$ through all its incident links. [^1] In order to model wireless radio communication, it is natural to assume that each agent in a dynamic network broadcasts its messages to all its neighbors (a message is received by anyone within communication range). The network’s anonymity prevents agents from specifying single destinations. By the end of round $t$, each agent reads all messages coming from its neighbors and updates its internal state according to a local algorithm $\mathcal{A}$. That is, $\mathcal{A}$ takes as input the current internal state of the agent and the multiset of incoming messages, and returns the new internal state of the agent. Note that $\mathcal{A}$ is deterministic and is the same for all agents.
The input of each agent also includes a leader flag. The agents whose leader flag is set are called leaders (or supervisors). We will use the symbol $\mathcal{N}_{n,\ell}$ to denote the set of all dynamic networks of $n$ agents with all possible input assignments, with the condition that exactly $\ell$ agents are leaders. We also define $\mathcal{N}_{n}=\bigcup_{\ell=0}^{n}\mathcal{N}_{n,\ell}$ as the set of networks of $n$ agents with an “unknown” number of leaders.
@@
Stabilization and termination. Each agent returns an output at the end of each round; the output is determined by its current internal state. A system is said to stabilize if the outputs of all its agents remain constant from a certain round onward; note that an agent’s internal state may still change even when its output is constant.
An agent may also decide to explicitly terminate and no longer update its internal state. An agent does so by setting a special termination flag in its state, which implies that no further state transitions are possible. Since the output of an agent only depends on its internal state, a terminated agent can no longer update its output, either. When all agents have terminated, the system is said to terminate, as well.
@@
Computation. We say that an algorithm $\mathcal{A}$ computes a function $F$ if, whenever the agents in the system are assigned inputs $\lambda(p_{1})$, $\lambda(p_{2})$, …, $\lambda(p_{n})$ and all agents execute the local algorithm $\mathcal{A}$ at every round, the system eventually stabilizes with each agent $p_{i}$ giving the desired output $F(p_{i},\lambda)$. A stronger notion of computation requires the system to not only stabilize but also to explicitly terminate with the correct output.
Formally, a function computed by a system of $n$ agents maps $n$-tuples of input values to $n$-tuples of output values. Writing such a function as $F(p_{i},\lambda)$ emphasizes that the output of an agent may depend on all agents’ inputs, as well as on the agent itself. That is, different agents may give different outputs.
The (worst-case) running time of an algorithm $\mathcal{A}$, as a function of $n$, is the maximum number of rounds it takes for the system to stabilize (and optionally terminate), taken across all possible dynamic networks of $n$ agents and all possible input assignments, assuming that all agents run $\mathcal{A}$.
Relationship between dynamic networks and Population Protocols. The dynamic network model considered in this paper bears some minor similarities with Population Protocols [4]. Indeed, both models describe systems where a set of entities exchange messages in a non-deterministic fashion and modify their internal states according to a deterministic local algorithm. However, a key difference is that in Population Protocols communication occurs only between two entities at a time; moreover, the two communicating entities have different roles: transmitter and receiver. This feature of Population Protocols automatically breaks symmetry between communicating entities, greatly simplifying problems such as Leader Election, Average Consensus, etc. It also eliminates all potential ambiguities arising from multiple entities in the same state sending equal messages to the same entity.
On the other hand, Population Protocols typically require the set of possible states to be independent of the number of entities $n$, while in our network model we allow agents to have an arbitrary amount of internal memory (although the algorithms presented in this paper require only polynomial memory in $n$).
Another feature of Population Protocols is global fairness, which roughly requires that the sequence of interactions between entities must allow any potentially reachable global state to eventually occur. This guarantees that the system can make progress despite an “adversarial” scheduler that might otherwise indefinitely block certain transitions. Our dynamic network model does not adopt the same notion of global fairness; however, a typical requirement is that the communication graph $G_{t}$ should be connected for every $t\geq 1$, or more generally the dynamic disconnectivity $\tau$ should be finite (see Section 2.3).
@@ 2.2 Classes of Functions
Multiset-based functions. Let $\mu_{\lambda}=\{(z_{1},m_{1}),(z_{2},m_{2}),\dots,(z_{k},m_{k})\}$ be the multiset of all agents’ inputs. That is, for all $1\leq i\leq k$, there are exactly $m_{i}$ agents $p_{j_{1}},p_{j_{2}},\dots,p_{j_{m_{i}}}$whose input is $z_{i}=\lambda(p_{j_{1}})=\lambda(p_{j_{2}})=\dots=\lambda(p_{j_{m_{i}}})$; note that $n=\sum_{i=1}^{k}m_{i}$. A function $F$ is multiset-based if it has the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}),\mu_{\lambda})$. That is, the output of each agent depends only on its own input and the multiset of all agents’ inputs. Multiset-based functions are also known as multi-aggregation or multi-aggregate functions [26, 27].
@@
Once an agent $p_{i}$ with input $\lambda(p_{i})$ has determined the multiset $\mu_{\lambda}$ of all agents’ inputs, it can immediately compute any desired function $\psi(\lambda(p_{i}),\mu_{\lambda})$ within the same local computation phase. ∎
@@
The problem of computing the mean of all input values is called Average Consensus  [8, 17, 18, 19, 30, 40, 52, 53, 56, 57, 58, 64, 68]. The frequency-based function $F_{IF}(p_{i},\lambda)=\frac{1}{n}\cdot\mu_{\lambda}$ is called Input Frequency function, and is complete for the class of frequency-based functions.
@@
Suppose that an agent $p_{i}$ has determined $\frac{1}{n}\cdot\mu_{\lambda}=\{(z_{1},m_{1}/n),(z_{2},m_{2}/n),\dots,(z_{k},m_{k}/n)\}$. Then it can immediately find the smallest integer $d>0$ such that $d\cdot(m_{i}/n)$ is an integer for all $1\leq i\leq k$. Note that $\frac{d}{n}\cdot\mu_{\lambda}$ is a multiset. Hence, in the same round, $p_{i}$ can compute any desired function $\psi(\lambda(p_{i}),\frac{d}{n}\cdot\mu_{\lambda})$, and thus any frequency-based function, by definition. ∎
@@ 2.3 Disconnected Networks
Although the network $G_{t}$ at each individual round may be disconnected, in this paper we assume dynamic networks to be $\tau$-union-connected. That is, there is a (smallest) parameter $\tau\geq 1$, called dynamic disconnectivity, such that the sum of any $\tau$ consecutive $G_{t}$’s is a connected multigraph (by definition, the sum of (multi)graphs is obtained by adding together their adjacency matrices). Thus, for all $i\geq 1$, the multigraph $\left(V,\bigcup_{t=i}^{i+\tau-1}E_{t}\right)$ is connected (by definition, the union of multisets is obtained by adding together the multiplicities of equal elements). We will use the symbol $\mathcal{N}^{\tau}_{n,\ell}$ (respectively, $\mathcal{N}^{\tau}_{n}$) to indicate the subset of $\mathcal{N}_{n,\ell}$ (respectively, $\mathcal{N}_{n}$) consisting of the networks whose dynamic disconnectivity is exactly $\tau$.
Our $\tau$-union-connected networks should not be confused with the $\tau$-interval-connected networks from [44]. In those networks, the intersection (as opposed to the union) of any $\tau$ consecutive $E_{t}$’s induces a connected (multi)graph. In particular, a $\tau$-interval-connected network is connected at every round, while a $\tau$-union-connected network may not be, unless $\tau=1$. Incidentally, a network is 1-interval-connected if and only if it is 1-union-connected.
Knowledge of $\tau$. We will argue that terminating computations are impossible without some a-priori knowledge about $\tau$. We say that a function $F(p_{i},\lambda)$ is trivial if and only if it is of the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}))$. That is, the output of any agent $p_{i}$ depends only on its own input $\lambda(p_{i})$ and not on the inputs of other agents; therefore, $F$ can be computed “locally” and does not require communications between agents.
@@
There is no algorithm that computes a non-trivial function with termination on all simple networks within $\bigcup_{\tau\geq 1}\bigcup_{n\geq 1}\mathcal{N}^{\tau}_{n}$.
@@
Assume for a contradiction that a non-trivial function $F(p_{i},\lambda)$ is computed with termination by an algorithm $\mathcal{A}$ in all simple networks in $\bigcup_{\tau\geq 1}\bigcup_{n\geq 1}\mathcal{N}^{\tau}_{n}$. Since $F$ is non-trivial, there is an input value $z$ and two distinct output values $y$ and $y^{\prime}$ with the following properties.
  • (i)
    If $p_{1}$ is the only agent in a network (i.e., $n=1$), and $p_{1}$ is assigned input $z$ and executes $\mathcal{A}$, it terminates in $t$ rounds with output $y$.
  • (ii)
    There exists a network size $\tilde{n}>1$ and an input assignment $\lambda$ with $\lambda(p_{1})=z$ such that, whenever the agents in a network of size $\tilde{n}$ are assigned input $\lambda$ and execute $\mathcal{A}$, the agent $p_{1}$ eventually terminates with output $y^{\prime}\neq y$.
Let us now consider a simple dynamic network of $\tilde{n}$ agents, where $p_{1}$ is kept disconnected from all other agents for the first $t$ rounds (hence $\tau>t$). Assign input $\lambda$ to the agents, and let them execute algorithm $\mathcal{A}$. Due to property (i), since $p_{1}$ is isolated for $t$ rounds and has no knowledge of $\tau$, it terminates in $t$ rounds with output $y$. This contradicts property (ii), which states that $p_{1}$ should terminate with output $y^{\prime}\neq y$. ∎
Proposition 2.3 can be strengthened and extended in various ways. For instance, the same result can be shown to hold even if all agents know $n$ in advance.
@@
For any fixed $n>1$, there is no algorithm that computes a non-trivial function with termination on all networks within $\bigcup_{\tau\geq 1}\mathcal{N}^{\tau}_{n}$.
@@
We follow the same proof structure and notation as Proposition 2.3, with $\tilde{n}=n$, except that property (i) is now different. Instead of having a network with a single agent, we have a network of $n$ agents arranged in a cycle, each of which has input $z$. By assumption, all agents terminate after $t$ rounds with output $y$.
Now consider a second network of $n$ agents, where $p_{1}$ is connected only to itself via a double self-loop for the first $t$ rounds. If we assign input $\lambda$ to the agents, with $\lambda(p_{1})=z$, then from the viewpoint of $p_{1}$ this network is indistinguishable from the previous one. Hence $p_{1}$ still terminates with output $y$, while the correct output is $y^{\prime}\neq y$. ∎
A limitation of Proposition 2.4 is that, unlike Proposition 2.3, it requires the presence of self-loops in the network. However, we can remove this requirement and prove a similar impossibility result for simple networks, at the cost of slightly narrowing the class of ”non-trivial functions” that cannot be computed.
Specifically, if we assume that $\lambda$ assigns the same input $z$ to at least three agents, we can arrange these agents in a proper cycle. From the perspective of the agents within the cycle, this configuration is indistinguishable from any cycle of arbitrary length in which all agents have input $z$. Therefore, the same impossibility result still applies, but is now limited to functions that are not constant over all input assignments where at least three agents share the same input.
Sufficiency of $\tau=1$. The following result will be used repeatedly to reduce computations on $\tau$-union connected networks to computations on 1-union-connected networks. Essentially, it states that any algorithm for 1-union-connected networks can be extended to $\tau$-union connected networks (where $\tau$ is known to all agents) at the cost of a factor $\tau$ in the running time, which is optimal.
Proposition 2.5.
If there is an algorithm $\mathcal{A}_{1}$ that computes a function $F$ in $\mathcal{N}^{1}_{n}$ within $f(n)$ rounds (with termination) for all $n\geq 1$, then for any fixed $\tau\geq 1$ there is an algorithm $\mathcal{A}_{\tau}$ that computes $F$ in $\mathcal{N}^{\tau}_{n}$ within $\tau\cdot f(n)$ rounds (with termination) for all $n\geq 1$. Moreover, if $A_{1}$ is optimal, then $\mathcal{A}_{\tau}$ is optimal.
Proof.
@@
Subdivide time into blocks of $\tau$ consecutive rounds, and consider the following algorithm $\mathcal{A}_{\tau}$. Each agent collects and stores all messages it receives within a same block, and updates its state all at once at the end of the block. This reduces any $\tau$-union-connected network $\mathcal{G}=((V,E_{t}))_{t\geq 1}$ to a 1-union-connected network $\mathcal{G}^{\prime}=((V,E^{\prime}_{t}))_{t\geq 1}$, where $E^{\prime}_{t}=\bigcup_{i=(t-1)\tau+1}^{t\tau}E_{i}$. Thus, if $F$ can be computed within $f(n)$ rounds in all 1-union-connected networks (which include $\mathcal{G}^{\prime}$), then $F$ can be computed within $\tau f(n)$ rounds in the original network $\mathcal{G}$.
@@
Conversely, consider a 1-union-connected network $\mathcal{G}$, and construct a $\tau$-union-connected network $\mathcal{G}^{\prime}$ by inserting $\tau-1$ empty rounds (i.e., rounds with no links at all) between every two consecutive rounds of $\mathcal{G}$. Since no information circulates during the empty rounds, if $F$ cannot be computed within $f(n)$ rounds in $\mathcal{G}$, then $F$ cannot be computed within $\tau f(n)$ rounds in $\mathcal{G}^{\prime}$ (recall that running times are measured in the worst case across all possible networks). Therefore, if $\mathcal{A}_{\tau}$ is not optimal, then $A_{1}$ is not optimal either. ∎
We point out that the argument in the proof of Proposition 2.5 is correct because algorithms are required to work for all multigraphs, as opposed to simple graphs only. Indeed, since an agent $p_{i}$ may receive multiple messages from the same agent $p_{j}$ within a same block, the resulting network $\mathcal{G}^{\prime}$ may have multiple links between $p_{i}$ and $p_{j}$ in a same round, even if $\mathcal{G}$ does not.
@@
Relationship between dynamic disconnectivity and the dynamic diameter. A concept closely related to the dynamic disconnectivity $\tau$ of a network is its dynamic diameter (or temporal diameter) $d$, which is defined as the maximum number of rounds it may take for information to travel from any agent to any other agent at any point in time [12, 46]. It is a simple observation that $\tau\leq d\leq\tau(n-1)$.
@@
We will use the symbol $\mathcal{N}^{[d]}_{n,\ell}$ to denote the subset of $\mathcal{N}_{n,\ell}$ consisting of the networks (with $n$ agents of which $\ell$ are leaders) whose dynamic diameter is exactly $d$. The square brackets around $d$ are used to avoid confusion with $\mathcal{N}^{\tau}_{n,\ell}$.